# The Dirichlet problem for hessian quotient type curvature equations in Minkowski space

### v1

@@ Abstract.
In this paper, we consider the Dirichlet problem for a class of prescribed Hessian quotient type curvature equations in Minkowski space. By establishing the a priori $C^{2}$ estimates, we prove the existence of smooth spacelike hypersurfaces with a class of prescribed Hessian quotient type curvature and homogeneous boundary data in the non-degenerate case.
@@ 1. Introduction
In this paper, we consider the Dirichlet problem for the following non-degenerate Hessian quotient type curvature equations in Minkowski space $\mathbb{R}^{n,1}$:
(1.1) $\frac{\sigma_{k}}{\sigma_{l}}(\lambda(\eta[M_{u}]))=\psi(X).$
When $k=n$ and $l=0$, this problem was studied by H. Jiao and the first author in [15] for general boundary data.
@@
with homogeneous boundary data in degenerate form. The current work aims to extend the results of [9] from Euclidean space to Minkowski space $\mathbb{R}^{n,1}$, whose induced metric is given by
@@
Due to the lack of gradient estimates, we restrict our analysis to the non-degenerate case $\psi=\psi(X)>0$, see Remark 3.2.
In contrast to Euclidean space, when differentiating equations twice in the Minkowski space, uncontrollable terms including the square of curvatures will appear. This is the reason why the Dirichlet problem for $k$-curvature equations
@@
remains unsolved for $33\leqslant k\leqslant n-$ 3, although it has been solved a long time ago for corresponding equations in Euclidean space, where
@@
where $\kappa[M_{u}]=(\kappa_{1},\cdots,\kappa_{n})$ denote the principal curvatures of $M_{u}$. The reader is referred to [4, 16, 17, 22, 18, 6, 7, 30] for the study of the Dirichlet problem for (1.2) and the quotient type curvature equation in Euclidean space.
In this paper, we are interested in the graph $M_{u}$ of a spacelike function $u$ defined on a bounded domain $\Omega\subset\mathbb{R}^{n}$, which is
$M_{u}:=\{(x,u(x)):x\in\Omega\subset\mathbb{R}^{n}\}.$
For the function $v(x)\in C^{2}(\Omega)$, we denote $v_{i}=\frac{\partial v}{\partial x_{i}}$, $v_{ij}=\frac{\partial^{2}v}{\partial x_{i}\partial x_{j}}$, and $Dv=(v_{1},\cdots,v_{n})$ and $D^{2}v=(v_{ij})$ represent the ordinary gradient and Hessian matrix.
In Minkowski space, we call a function $u$ spacelike if
$\sup_{\overline{\Omega}}|Du|<1.$
@@
Define the $(0,2)$-tensor field $\eta$ on $M$ by
$\eta=Hg-h,$
where $H(X)$ is the mean curvature of $M_{u}$ at $X=(x,u(x))\in M_{u}$, $g=\{g_{ij}\}$, $h=\{h_{ij}\}$. $\lambda(\eta[M_{u}])=(\lambda_{1},\cdots,\lambda_{n})$ denote the eigenvalues of $\eta$ with respect to $g$ on $M_{u}$.
In this paper, we consider the existence of smooth spacelike graphic hypersurface satisfying the Dirichlet problem
(1.3) $\frac{\sigma_{k}}{\sigma_{l}}(\lambda(\eta[M_{u}]))&=\psi(x,u)&\;\;~{}\mbox{ in }\Omega,\\ u&=0&\;\;~{}\mbox{ on }\partial\Omega,\end{aligned}\right.$
@@
We first introduce the definition of admissible set of equation (1.1).
@@
A $C^{2}$ regular spacelike hypersurface $M_{u}$ is called $(\eta,k)$-convex if the principal curvatures $\kappa=(\kappa_{1},\ldots,\kappa_{n})\in\widetilde{\Gamma}_{k}$ at each $X\in M_{u}$, where $\widetilde{\Gamma}_{k}$ is the symmetric cone defined by
@@
We also need some geometric conditions on $\Omega$ similar to those in [17] when consider the Dirichlet boundary value problem.
Definition 1.2.
If $k<n$, a bounded domain $\Omega\subset\mathbb{R}^{n}$ is called $(\eta,k)$-convex, if there exists a positive constant $K$ such that for each $x\in\partial\Omega$,
$(\kappa^{b}_{1}(x),\ldots,\kappa^{b}_{n-1}(x),K)\in\widetilde{\Gamma}_{k},$
where $\kappa^{b}_{1}(x),\ldots,\kappa^{b}_{n-1}(x)$ are the principal curvatures of $\partial\Omega$ at $x$.
Jiao and Sun [20] considered the following degenerate curvature equation
@@
in Euclidean space with constant boundary condition. They derived the existence of $C^{1,1}$ regular graphical $(\eta,n)$-convex hypersurfaces. Motivated by [20], Chen-Tu-Xiang [9] extended this work to a degenerate Hessian quotient type curvature equation (1.1) with $0\leqslant l<k<n$, they proved the existence and uniqueness of $C^{1,1}$ regular graphical hypersurfaces in Euclidean space under the condition $\psi^{\frac{1}{k-1}}(x,u(x),\nu(x))\in C^{1,1}(\Omega\times\mathbb{R}\times S^{n})$. Building on [9], we generalize their results to Minkowski space and establish the following theorem:
@@
Let $k\geqslant 2,~{}0\leqslant l<k<n$. Assume that bounded domain $\Omega\subset\mathbb{R}^{n}$ is $(\eta,k)$-convex with smooth boundary $\partial\Omega$. Suppose $\psi=\psi(x,z)\in C^{\infty}(\overline{\Omega}\times\mathbb{R})>0$, $\psi_{z}\geqslant 0$. In addition, assume that there exists an admissible subsolution $\underline{u}\in C^{2}(\overline{\Omega})$ satisfying $\kappa[M_{\underline{u}}]\in\widetilde{\Gamma}_{k}$ and
[…]
@@
To conclude the introduction, we briefly mention some related studies. If $\lambda(\eta[M_{u}])$ is replaced by the principal curvature $\kappa[M_{u}]$ of $M_{u}$ and $l=0$, equation (1.1) becomes the classical $k$-curvature equations (1.2). The mean curvature, scalar curvature, and Gauss curvature correspond respectively to $k=1,2,n$ in the formula (1.2).
The Dirichlet problem for prescribed mean curvature in Minkowski space was first solved by Bartnik and Simon [1], while Delanoè [10] studied the Dirichlet problem of the prescribed Gauss curvature equation (see also [11]). Bayard [2] completely solved the solvability of the Dirichlet problem of the prescribed scalar curvature equation in four-dimensional space under the condition of strict convexity, and Urbas [28] generalized the result to the general dimension $n$. Ren and Wang [24, 25] solved the case $k=n-1$ and $k=n-2$. For $33\leqslant k\leqslant n-$ 3, this is still an open problem. The reason is the lack of a priori $C^{2}$ estimates. For this problem, Guan-Ren-Wang [14] given a second order estimates for admissible solutions of $k$-curvature equations (1.2) whose graphs are $(k+1)$-convex, and Wang [29] established the estimates for admissible solutions whose graphs are semi-convex.
@@
Liu-Mao-Zhao [23] established Pogorelov type estimates of $k$-convex solutions with the vanishing Dirichlet boundary condition in Lorentz-Minkowski space $\mathbb{R}^{n+1}$. Chen-Tu-Xiang [8] considered the Dirichlet problem for the above equation on compact Riemannian Manifolds and obtained the existence and uniqueness.
The rest of this paper is organized as follows. In Section 2, we provided some preliminaries. In Section 3, The $C^{1}$ estimates are established. Section 4 is devoted to estimates for second order derivatives.
@@ 2. Preliminaries
[…]
@@
Define the martix $A=\{a_{ij}\}$, where
(2.2) $a_{ij}=\frac{1}{w}\gamma^{ik}u_{kl}\gamma^{lj},$
and $w=\sqrt{1-|Du|^{2}}$, $\gamma^{ik}=\delta_{ik}+\frac{u_{i}u_{k}}{w(1+w)}.$ Note that $\{\gamma^{ij}\}$ is invertible and its inverse is $\{\gamma_{ij}\}=\{\delta_{ij}-\frac{u_{i}u_{j}}{1+w}\}$, which is the square root of $\{g_{ij}\}.$
@@
$\lambda(\{a_{ij}\})=\lambda(\{g^{ij}\}\{h_{ij}\})=\lambda\big{(}\frac{1}{w}\left(I+\frac{Du\otimes Du}{w^{2}}\right)D^{2}u\big{)}.$
@@
In the final part of this section, we list some algebraic equalities and inequalities of $\sigma_{k}$. Denote $\sigma_{m;i_{1},\cdots,i_{k}}(\kappa)=\sigma_{m}(\kappa)|_{\kappa_{i_{1}}=\cdots=\kappa_{i_{k}}=0}$ for integer $1\leqslant i_{1},\cdots,i_{k}\leqslant n,1\leqslant m\leqslant n$ and $n-k\leqslant m$. Then we have:
@@
For $\kappa\in\Gamma_{k}$ and $n\geqslant k>l\geqslant 0$, $n\geqslant r>s\geqslant 0$, $k\geqslant r$, $l\geqslant s$, we have
$\Bigg{[}\frac{{\sigma_{k}(\kappa)}/{C_{n}^{k}}}{{\sigma_{l}(\kappa)}/{C_{n}^{l}}}\Bigg{]}^{\frac{1}{k-l}}\leqslant\Bigg{[}\frac{{\sigma_{r}(\kappa)}/{C_{n}^{r}}}{{\sigma_{s}(\kappa)}/{C_{n}^{s}}}\Bigg{]}^{\frac{1}{r-s}}.$
@@
Next, we introduce a new cone $\widetilde{\Gamma}_{k}$ defined by (1.4) and list some important properties, more details can refer to [5].
@@
(2.4) $\Gamma_{1}=\widetilde{\Gamma}_{1}\supset\widetilde{\Gamma}_{2}\supset\cdots\supset\widetilde{\Gamma}_{n}\supset\Gamma_{2}.$
@@ 3. $C^{1}$ estimates
In this section, we establish the $C^{1}$ estimates for the admissible solution to (1.3). Simliar as [15], we prove the admissible solution $u$ is spacelike, which means there exists a constant $10<\theta_{0}<$ 1 such that
@@
Since $M_{u}$ is $(\eta,k)$-convex, then $\sigma_{1}(\lambda(\eta[M_{u}]))>0$, hence $H[M_{u}]>0$. By proposition (2.1), we can find a spacelike supersolution $\overline{u}\in C^{\infty}(\overline{\Omega})$ of (1.3). Then combining with $\psi_{u}\geqslant 0$, the subsolution condition (1.5) and the maximum principle, we can find that
@@
for some constant $10<\theta<$ 1 since $\underline{u}$ and $\overline{u}$ are both spacelike. Next we establish the global gradient estimates (3.1), that is, we prove that the function
@@
(3.4) $\sup_{\overline{\Omega}}\tilde{w}\leqslant C(1+\sup_{\partial\Omega}\tilde{w}),$
@@
where $\gamma^{is}:=\delta_{is}+\frac{u_{i}u_{s}}{w(1+w)}$ and ~$\tilde{\partial}_{s}:=\epsilon_{s}+u_{s}\epsilon_{n+1}$. It’s clear that $\{e_{1},e_{2},\cdots,e_{n}\}$ be an orthonormal frame on $TM_{u}$ around $X_{0}=(x_{0},u(x_{0}))$ such that, at $x_{0}$,
(3.5) $\nabla_{1}u=\frac{|Du|}{w}=|\nabla u|,\nabla_{i}u=u_{i}=0,\text{ for }i\geqslant 2.$
@@
(3.7) $0\geqslant F^{ii}\nabla_{ii}Q=\frac{F^{ii}\nabla_{i}h_{i1}\nabla_{1}u+F^{ii}h_{il}\nabla_{il}u}{\tilde{w}}-\frac{F^{ii}\left(h_{i1}\nabla_{1}u\right)^{2}}{\tilde{w}^{2}}+BF^{ii}\nabla_{ii}u,$
@@
Remark 3.2.
Although Chen-Tu-Xiang [9] has given the gradient estimate for admissible solutions of the equation
$\frac{\sigma_{k}}{\sigma_{l}}(\lambda(\eta[M_{u}]))=\psi(x,u,\nu)\geqslant 0$
in Euclidean space, gradient estimates in Minkowski space remain a challenge for degenerate problems due to the metric. In addition, no satisfactory approach has been developed to control the term $\nu$, even in non-degenerate cases.

### v2

@@ Abstract.
In this paper, we consider the Dirichlet problem for a class of prescribed Hessian quotient type curvature equations with homogeneous boundary data in Minkowski space. By establishing the a priori $C^{2}$ estimates, we obtain the existence result in the non-degenerate case.
@@ 1. Introduction
In Minkowski space $\mathbb{R}^{n,1}$, a function $u$ defined on a bounded domain $\Omega\subset\mathbb{R}^{n}$ is called spacelike if
$\sup_{\overline{\Omega}}|Du|<1.$
In this paper, we are interested in the prescribed curvature problems for graph $M_{u}$ of a spacelike function $u$, which is
$M_{u}:=\{(x,u(x)):x\in\Omega\subset\mathbb{R}^{n}\}.$
We shall consider the Dirichlet problem for the following non-degenerate Hessian quotient type curvature equations:
(1.1) $\frac{\sigma_{k}}{\sigma_{l}}(\lambda(\eta[M_{u}]))=\psi(X)\;\mbox{in}\;\Omega$
with homogeneous boundary data. Here $X=(x,u(x))$ is the position vector, non-degenerate means $\psi(X)>0$, $\sigma_{k}$ is the $k$-th elementary symmetric function, $\lambda(\eta[M_{u}])=(\lambda_{1},\cdots,\lambda_{n})$ denotes the eigenvalues of $\eta$ with respect to $g$ on $M_{u}$. $\eta$ is the $(0,2)$-tensor field defined on $M$ by
$\eta=Hg-h,$
where $H(X)$ is the mean curvature of $M_{u}$ at $X\in M_{u}$, $g=\{g_{ij}\}$ is the induced Riemannian metric and $h=\{h_{ij}\}$ is the second fundamental form.
When $k=n$ and $l=0$, this problem was studied by Jiao and the first author in [16] for general boundary data.
@@
with homogeneous boundary data in the degenerate case. The current work aims to extend the results of [10] from Euclidean space to Minkowski space $\mathbb{R}^{n,1}$, whose induced metric is given by
@@
In contrast to the Euclidean case, when differentiating equations twice in the Minkowski space, uncontrollable terms including the square of curvatures will appear. Hence the Dirichlet problem for $k$-curvature equations
@@
remains unsolved for $3\leqslant k\leqslant n-3$, although it has been solved a long time ago for its Euclidean analogue, where
@@
$\kappa[M_{u}]=(\kappa_{1},\cdots,\kappa_{n})$ denotes the principal curvatures of $M_{u}$. The reader is referred to [5, 7, 8, 15, 17, 18, 19, 23, 25, 33] for the study of the Dirichlet problem for (1.2) and the quotient type curvature equation in Euclidean space.
For the function $v(x)\in C^{2}(\Omega)$, we denote $v_{i}=\frac{\partial v}{\partial x_{i}}$, $v_{ij}=\frac{\partial^{2}v}{\partial x_{i}\partial x_{j}}$, and use $Dv=(v_{1},\cdots,v_{n})$ and $D^{2}v=\{v_{ij}\}$ to represent the ordinary gradient and Hessian matrix respectively.
@@
In this paper, we consider the existence of smooth spacelike graphic hypersurfaces satisfying the Dirichlet problem
(1.3) $\left\{\begin{aligned} \frac{\sigma_{k}}{\sigma_{l}}(\lambda(\eta[M_{u}]))&=\psi(x,u)&\;\;~\mbox{ in }\Omega,\\ u&=0&\;\;~\mbox{ on }\partial\Omega,\end{aligned}\right$.
@@
Define the Gårding’s cone
We now introduce the definition of admissible set of equation (1.1).
@@
A $C^{2}$ regular spacelike hypersurface $M$ is called $(\eta,k)$-convex if the principal curvatures $\kappa=(\kappa_{1},\ldots,\kappa_{n})\in\widetilde{\Gamma}_{k}$ at each $X\in M$, where $\widetilde{\Gamma}_{k}$ is the symmetric cone defined by
@@
Jiao-Sun [21] considered the following degenerate curvature equation
@@
in Euclidean space with constant boundary condition. They derived the existence of $C^{1,1}$ regular graphical $(\eta,n)$-convex hypersurfaces. Motivated by [21], Chen-Tu-Xiang [10] extended this work to a degenerate Hessian quotient type curvature equation (1.1) with $0\leqslant l<k<n$ and proved the existence and uniqueness of $C^{1,1}$ regular graphical hypersurfaces in Euclidean space under the condition $\psi^{\frac{1}{k-1}}(x,u(x),\nu(x))\in C^{1,1}(\Omega\times\mathbb{R}\times S^{n})$. Building on [10], we generalize their results to Minkowski space and establish the following theorem:
@@
Let $k\geqslant 2,~0\leqslant l<k<n$. Assume that $\Omega\subset\mathbb{R}^{n}$ is bounded domain with smooth boundary $\partial\Omega$. Suppose $\psi=\psi(x,z)\in C^{\infty}(\overline{\Omega}\times\mathbb{R})>0$, $\psi_{z}\geqslant 0$. In addition, assume that there exists an admissible subsolution $\underline{u}\in C^{2}(\overline{\Omega})$ satisfying $\kappa[M_{\underline{u}}]\in\widetilde{\Gamma}_{k}$ and
(1.5) $\left\{\begin{aligned} &\frac{\sigma_{k}}{\sigma_{l}}(\lambda(\eta[M_{\underline{u}}]))\geqslant\psi^{0}(x,\underline{u})\geqslant\psi(x,u),&&in~\Omega,\\ &\underline{u}=0,&&on~\partial\Omega\end{aligned}\right$.
@@
Generally speaking, we need some geometric conditions on $\Omega$ as in [18] to deal with equation (1.1). Such convexity-type conditions are typically used to derive subsolutions or construct barriers in second order estimates on the boundary for Hessian equation and curvature equations in the form
$F(D^{2}u)=f(\lambda(D^{2}u))=\psi\;\mbox{and}\;F(\kappa(M))=\psi.$
For equations like (1.1), the correspond condition is: there exists a positive constant $K$ such that for each $x\in\partial\Omega$,
(1.6) $(\kappa_{1}^{\prime}(x),\dots,\kappa_{n-1}^{\prime}(x),K)\in\widetilde{\Gamma}_{k},$
where $\kappa_{1}^{\prime}(x),\dots,\kappa_{n-1}^{\prime}(x)$ are the principal curvatures of $\partial\Omega$ at $x$. However, we notice that for equation (1.1), when $k<n$,
$\sigma_{k}(\lambda(\kappa_{1}^{\prime},\cdots\kappa_{n-1}^{\prime},K))=C_{n-1}^{k}K^{k}+\mbox{bounded terms},$
where $\lambda=(\lambda_{1},\cdots,\lambda_{n})$ is defined as in (1.4). It follows that, for any $K$ sufficiently large, $\sigma_{k}(\lambda(\kappa_{1}^{\prime},\cdots\kappa_{n-1}^{\prime},K))$ is always positive and thus,
$(\kappa_{1}^{\prime}(x),\dots,\kappa_{n-1}^{\prime}(x),K)\in\widetilde{\Gamma}_{q},\;q=1,2,\ldots,n-1.$
In fact, since the positive $\kappa_{i}$ axes belong to $\tilde{\Gamma}_{k}$ for $k<n$, it is the cone of type 2 (see [4] for the definition) and condition (1.6) holds automatically. Therefore, the geometric restrictions on $\Omega$ can be dropped in Theorem 1.2.
To conclude the introduction, we briefly mention some related studies. If $\lambda(\eta[M_{u}])$ is replaced by the principal curvature $\kappa$ of $M_{u}$ and $l=0$, equation (1.1) becomes the classical $k$-curvature equation (1.2). The mean curvature, scalar curvature and Gauss curvature equations correspond respectively to $k=1,2,$ and $n$ in the formula (1.2).
The Dirichlet problem for the prescribed mean curvature equation in Minkowski space was first solved by Bartnik-Simon [1], while Delanoè [11] studied the Dirichlet problem of the prescribed Gauss curvature equation (see also [12]). Bayard [2] solved the Dirichlet problem for the prescribed scalar curvature equation in dimension four under the condition of strict convexity, and Urbas [30] generalized the result to general dimension. Ren-Wang [26, 27] solved the case k=n1k=n-1 and k=n2k=n-2. For $3\leqslant k\leqslant n-3$, the problem remains open, due to the lack of a priori $C^{2}$ estimates. Wang [32] obtained the curvature estimates for (k+1)-convex solutions of $k$-curvature equations (1.2) in Minkowski space, while Wang [31] established the estimates under semi-convex condition.
@@
Liu-Mao-Zhao [24] established Pogorelov type estimates of $k$-convex solutions with homogeneous boundary data in Lorentz-Minkowski space $\mathbb{R}^{n+1}$. Chen-Tu-Xiang [9] considered the Dirichlet problem on compact Riemannian manifolds and obtained the existence result.
The rest of this paper is organized as follows. In Section 2, we provide some preliminaries. The $C^{1}$ estimate is estabished in Section 3. We deal with the boundary second order derivative estimates in Section 4. Finally, the proof of Theorem 1.2 is completed in Section 5.
@@ 2. Preliminaries
Then $u(x)=-\langle X,\epsilon_{n+1}\rangle$, where $\langle\cdot,\cdot\rangle$ is the Minkowski inner product.
@@
Define
(2.2) $A=\{a_{ij}\}=\Big\{\frac{1}{w}\gamma^{ik}u_{kl}\gamma^{lj}\Big\},$
where $w=\sqrt{1-|Du|^{2}}$ and $\gamma^{ik}=\delta_{ik}+\frac{u_{i}u_{k}}{w(1+w)}.$ Note that $\{\gamma^{ij}\}$ is invertible and its inverse is $\{\gamma_{ij}\}=\{\delta_{ij}-\frac{u_{i}u_{j}}{1+w}\}$, which is the square root of $\{g_{ij}\}.$
@@
(2.3) $\kappa(\{a_{ij}\})=\kappa(\{g^{ij}\}\{h_{ij}\})=\kappa\big(\frac{1}{w}\left(I+\frac{Du\otimes Du}{w^{2}}\right)D^{2}u\big).$
@@
In the final part of this section, we list some algebraic equalities and inequalities of $\sigma_{k}$. Denote $\sigma_{m;i_{1},\cdots,i_{k}}(\kappa)=\sigma_{m}(\kappa)|_{\kappa_{i_{1}}=\cdots=\kappa_{i_{k}}=0}$ for integer $1\leqslant i_{1},\cdots,i_{k}\leqslant n,1\leqslant m\leqslant n$ and $n-k\leqslant m$. Then we have,
@@
For $\lambda\in\Gamma_{k}$ and $n\geqslant k>l\geqslant 0$, $n\geqslant r>s\geqslant 0$, $k\geqslant r$, $l\geqslant s$, we have
$\Bigg[\frac{{\sigma_{k}(\lambda)}/{C_{n}^{k}}}{{\sigma_{l}(\lambda)}/{C_{n}^{l}}}\Bigg]^{\frac{1}{k-l}}\leqslant\Bigg[\frac{{\sigma_{r}(\lambda)}/{C_{n}^{r}}}{{\sigma_{s}(\lambda)}/{C_{n}^{s}}}\Bigg]^{\frac{1}{r-s}}.$
@@
Next, we list some important properties cone $\widetilde{\Gamma}_{k}$ defined by (1.4), more details can refer to [6].
@@
(2.5) $\Gamma_{1}=\widetilde{\Gamma}_{1}\supset\widetilde{\Gamma}_{2}\supset\cdots\supset\widetilde{\Gamma}_{n}\supset\Gamma_{2}\supset\Gamma_{3}\cdots\supset\Gamma_{n}.$
@@ 3. $C^{1}$ estimates
Next, we establish the $C^{1}$ estimates for the admissible solution to (1.3). Simliar as [16], we prove the admissible solution $u$ is spacelike, which means there exists a constant $0<\theta_{0}<1$ such that
@@
This global gradient estimates will give an upper bound of $\tilde{w}$, which ensures the uniform ellipticity of $\frac{\sigma_{k}}{\sigma_{l}}(\lambda(\eta))$ if $u||_{C^{2}(\overline{\Omega})}<C$.
Since $M_{u}$ is $(\eta,k)$-convex, we have $\sigma_{1}(\lambda(\eta[M_{u}]))>0$, which means $\sigma_{1}(\kappa(M_{u}))=H[M_{u}]>0$. Obviously, $\bar{u}=0$ is a supersolution of Dirichlet problem (1.3). Combining with $\psi_{u}\geqslant 0$, the subsolution condition (1.5) and the maximum principle, we can find that
$\underline{u}\leqslant u\leqslant\overline{u}\text{ in }\Omega\text{ and }\overline{u}=u=\underline{u}\text{ on }\partial{\Omega},$
which implies
$\frac{\partial\underline{u}}{\partial\gamma}\leqslant\frac{\partial u}{\partial\gamma}\leqslant\frac{\partial\overline{u}}{\partial\gamma},$
where $\gamma$ is the interior unit normal to $\partial\Omega$. Then we have
@@
for some constant $0<\theta<1$ since $\underline{u}$ and $\overline{u}$ are both spacelike.
Next, we prove an upper bound for
@@
[…]
@@
where $\gamma^{is}:=\delta_{is}+\frac{u_{i}u_{s}}{w(1+w)}$ and $\tilde{\partial}_{s}:=\epsilon_{s}+u_{s}\epsilon_{n+1}$. It’s clear that $\{e_{1},e_{2},\cdots,e_{n}\}$ is an orthonormal frame on $TM_{u}$ around $X_{0}=(x_{0},u(x_{0}))$ such that, at $x_{0}$,
(3.5) $\nabla_{1}u=\frac{|Du|}{w}=|\nabla u|,\nabla_{i}u=\frac{u_{i}}{w}=0,\text{ for }i\geqslant 2.$
@@
(3.7) $0\geqslant F^{ij}\nabla_{ij}Q=\frac{F^{ij}\nabla_{j}h_{i1}\nabla_{1}u+F^{ij}h_{il}\nabla_{jl}u}{\tilde{w}}-\frac{F^{ij}(h_{i1}\nabla_{1}u)(h_{j1}\nabla_{1}u)}{\tilde{w}^{2}}+BF^{ij}\nabla_{ij}u,$